# First Heuristic Then Rational: Dynamic Use of Heuristics in Language Model Reasoning

### v1

@@ Appendix G Effect of few-shot examples
We investigate whether the model’s heuristic is triggered by the few-shot examples. Specifically, we replace the few-shot examples in the following ways to study the relationship between the model’s heuristic and its inputs:
@@
3. We change the few-shot examples to induce Neg. (as shown in Table 13), and investigate if there’s a decrease in reasoning frequency with the use of distractors in the Neg. dataset compared to Table 1.
We measure the frequency of selecting $\tilde{p}$ in the settings of  § 4.1. The results are presented in Table 8. As shown in Tables 18, although the few-shot examples fed into the models such as GPT-3.5, GPT-4, and PaLM2 was changed, there was no significant change in reasoning frequency as described. This suggests that the model’s heuristic does not merely mimic the examples provided as input. On the other hand, the Llama2 model was more prone to being misled by changes in input, and smaller models demonstrated a reduced capacity to reach the correct answers directory.
Answer the context question according to the following example.
@@
Context: Alice has 92 more bananas than Quentin’s mother. Victor has 10 less bananas than Walter. Xavier has 59 more bananas than Sybil. Yvonne has 79 more bananas than Sybil. Judy has 23 more bananas than Alice. Dave has 60 more bananas than Victor. Quentin has 35 less bananas than Quentin’s father. Heidi has 95 more bananas than Victor. Ursula doesn’t have 32 more bananas than Quentin’s father. Larry has 17 less bananas than Alice. Zoe has 58 less bananas than Yvonne. Ivan has 43 less bananas than Yvonne. Walter has 43 less bananas than Quentin’s mother. Nancy has 34 bananas. Grace has 41 more bananas than Xavier. Quentin’s mother has 55 less bananas than Nancy. Sybil has 3 less bananas than Nancy. Quentin’s father has 50 more bananas than Walter. Trent has 33 less bananas than Xavier.
@@
Nancy has 34 bananas, and Quentin’s mother has 55 less bananas than Nancy. So, Quentin’s mother has 34-55=-21 bananas.
Quentin’s mother has -21 bananas, and Walter has 43 less bananas than Quentin’s mother. So, Walter has -21-43=-64 bananas.
@@
Quentin’s father has -14 bananas, and Quentin has 35 less bananas than Quentin’s father. So, Quentin has -14-35=-49 bananas.
@@
Context: Yvonne’s father has 10 more apples than Yvonne’s son. Eve has 2 apples. Yvonne’s son has 3 more apples than Eve. Quentin has 3 more apples than Yvonne. Yvonne has 3 less apples than Yvonne’s father. Alice has 3 more apples than Grace. Trent has 34 more apples than Yvonne’s father. Ivan has 3 apples. Ursula has 3 more apples than Yvonne’s father. Grace has 3 apples. Xavier has 3 more apples than Ivan.
@@
Yvonne’s father has 15 apples, and Yvonne has 3 less apples than Yvonne’s father. So, Yvonne has 15-3=12 apples.
@@
Context: Kevin’s friend has 33 less grapes than Rob. Ivan has 43 more grapes than Victor. Victor has 33 less grapes than Kevin’s friend. Ursula has 75 less grapes than Zoe. Alice has 11 more grapes than Eve. Dave has 11 more grapes than Eve. Olivia has 29 more grapes than Kevin’s friend. Mallory has 97 more grapes than Olivia. Judy has 78 more grapes than Olivia. Rob has 55 grapes. Frank has 70 less grapes than Heidi. Eve has 84 less grapes than Kevin’s neighborhood. Xavier has 36 more grapes than Heidi. Kevin’s neighborhood has 55 less grapes than Kevin’s friend. Kevin has 43 less grapes than Kevin’s mother. Heidi has 61 less grapes than Kevin’s friend. Kevin’s mother has 88 more grapes than Kevin’s neighborhood. Kevin’s friend has 40 more grapes than Rob. Walter has 38 more grapes than Victor.
@@
Kevin’s friend has 95 grapes, and Kevin’s neighborhood has 55 less grapes than Kevin’s friend. So, Kevin’s neighborhood has 95-55=40 grapes.
@@
Kevin’s mother has 128 grapes, and Kevin has 43 less grapes than Kevin’s mother. So, Kevin has 128-43=85 grapes.
The final answer is 128.
@@
Answer the context question according to the following example.
@@
Context: Nancy has 34 bananas. Mallory has 55 less bananas than Nancy. Walter has 43 less bananas than Mallory. Peggy has 50 more bananas than Walter. Quentin has 35 less bananas than Peggy. Alice has 92 more bananas than Mallory. Victor has 10 less bananas than Walter. Xavier has 59 more bananas than Sybil. Yvonne has 79 more bananas than Sybil. Judy has 23 more bananas than Alice. Dave has 60 more bananas than Victor. Heidi has 95 more bananas than Victor. Ursula doesn’t have 32 more bananas than Peggy. Larry has 17 less bananas than Alice. Zoe has 58 less bananas than Yvonne. Ivan has 43 less bananas than Yvonne. Grace has 41 more bananas than Xavier. Sybil has 3 less bananas than Nancy. Trent has 33 less bananas than Xavier.
@@
Nancy has 34 bananas, and Mallory has 55 less bananas than Nancy. So, Mallory has 34-55=-21 bananas.
Mallory has -21 bananas, and Walter has 43 less bananas than Mallory. So, Walter has -21-43=-64 bananas.
@@
Peggy has -14 bananas, and Quentin has 35 less bananas than Peggy. So, Quentin has -14-35=-49 bananas.
@@
Context: Eve has 2 apples. Yvonne’s son has 3 more apples than Eve. Zoe has 10 more apples than Yvonne’s son. Yvonne has 3 less apples than Zoe. Alice has 3 more apples than Grace. Quentin has 3 more apples than Yvonne. Trent has 34 more apples than Zoe. Ivan has 3 apples. Ursula has 3 more apples than Zoe. Grace has 3 apples. Xavier has 3 more apples than Ivan.
@@
Zoe has 15 apples, and Yvonne has 3 less apples than Zoe. So, Yvonne has 15-3=12 apples.
@@
Context: Rob has 55 grapes. Trent has 40 more grapes than Rob. Sybil has 55 less grapes than Trent. Zoe has 88 more grapes than Sybil. Kevin has 43 less grapes than Zoe. Kevin’s friend has 33 less grapes than Rob. Ivan has 43 more grapes than Victor. Victor has 33 less grapes than Kevin’s friend. Ursula has 75 less grapes than Zoe. Alice has 11 more grapes than Eve. Dave has 11 more grapes than Eve. Olivia has 29 more grapes than Kevin’s friend. Mallory has 97 more grapes than Olivia. Judy has 78 more grapes than Olivia. Frank has 70 less grapes than Heidi. Eve has 84 less grapes than Sybil. Xavier has 36 more grapes than Heidi. Heidi has 61 less grapes than Trent. Walter has 38 more grapes than Victor.
@@
Trent has 95 grapes, and Sybil has 55 less grapes than Trent. So, Sybil has 95-55=40 grapes.
@@
Zoe has 128 grapes, and Kevin has 43 less grapes than Zoe. So, Kevin has 128-43=85 grapes.
The final answer is 128.
@@
Answer the context question according to the following example.
@@
Context: Alice has 92 more bananas than Mallory. Victor has 10 less bananas than Walter. Xavier has 59 more bananas than Sybil. Yvonne doesn’t have 79 more bananas than Sybil. Judy doesn’t have 23 more bananas than Alice. Dave has 60 more bananas than Victor. Quentin has 35 less bananas than Peggy. Heidi has 95 more bananas than Victor. Ursula doesn’t have 32 more bananas than Peggy. Larry doesn’t have 17 less bananas than Alice. Zoe has 58 less bananas than Yvonne. Ivan has 43 less bananas than Yvonne. Walter has 43 less bananas than Mallory. Nancy has 34 bananas. Grace doesn’t have 41 more bananas than Xavier. Mallory has 55 less bananas than Nancy. Sybil doesn’t have 3 less bananas than Nancy. Peggy has 50 more bananas than Walter. Trent doesn’t have 33 less bananas than Xavier.
@@
Nancy has 34 bananas, and Mallory has 55 less bananas than Nancy. So, Mallory has 34-55=-21 bananas.
Mallory has -21 bananas, and Walter has 43 less bananas than Mallory. So, Walter has -21-43=-64 bananas.
@@
Peggy has -14 bananas, and Quentin has 35 less bananas than Peggy. So, Quentin has -14-35=-49 bananas.
@@
Context: Zoe has 10 more apples than Yvonne’s son. Eve has 2 apples. Yvonne’s son has 3 more apples than Eve. Quentin has 3 more apples than Yvonne. Yvonne has 3 less apples than Zoe. Alice has 3 more apples than Grace. Trent has 34 more apples than Zoe. Ivan has 3 apples. Ursula has 3 more apples than Zoe. Grace has 3 apples. Xavier doesn’t have 3 more apples than Ivan.
@@
Zoe has 15 apples, and Yvonne has 3 less apples than Zoe. So, Yvonne has 15-3=12 apples.
@@
Context: Kevin’s friend has 33 less grapes than Rob. Ivan doesn’t have 43 more grapes than Victor. Victor doesn’t have 33 less grapes than Kevin’s friend. Ursula has 75 less grapes than Zoe. Alice has 11 more grapes than Eve. Dave has 11 more grapes than Eve. Olivia doesn’t have 29 more grapes than Kevin’s friend. Mallory has 97 more grapes than Olivia. Judy doesn’t have 78 more grapes than Olivia. Rob has 55 grapes. Frank has 70 less grapes than Heidi. Eve has 84 less grapes than Sybil. Xavier doesn’t 36 more grapes than Heidi. Sybil has 55 less grapes than Trent. Kevin has 43 less grapes than Zoe. Heidi has 61 less grapes than Trent. Zoe has 88 more grapes than Sybil. Trent has 40 more grapes than Rob. Walter has 38 more grapes than Victor.
@@
Trent has 95 grapes, and Sybil has 55 less grapes than Trent. So, Sybil has 95-55=40 grapes.
@@
Zoe has 128 grapes, and Kevin has 43 less grapes than Zoe. So, Kevin has 128-43=85 grapes.
The final answer is 128.

### v2

@@ Appendix G Effect of few-shot examples
We investigate whether the few-shot examples trigger the model’s heuristic. Specifically, we replace the few-shot examples in the following ways to study the relationship between the model’s heuristic and its inputs:
@@
3. We change the few-shot examples to induce Negative (as shown in Table 13), and investigate if there’s a decrease in reasoning frequency with the use of distractors in the Negative dataset compared to Table 1.
We measure the frequency of selecting $\tilde{p}$ in § 4.1. The results are presented in Table 8. As shown in Tables 18, although the few-shot examples fed into the models such as GPT-3.5, GPT-4, and PaLM2 was changed, there was no significant change in reasoning frequency as described. This suggests that the model’s heuristic does not merely mimic the examples provided as input. On the other hand, the Llama2 model was more prone to being misled by changes in input, and smaller models demonstrated a reduced capacity to reach the correct answers.
Appendix H Effects of increasing the number of distructors
We investigate whether the same results could be obtained when the number of distractors that do not include heuristics increases. Specifically, we prepare the same settings as in § 4.2 and add distractors that do not include heuristics. We define the distractors that include heuristics as $\tilde{p}_{t,\text{heuristic}}$ and those that do not as $\tilde{p}_{t,\text{non-heuristic}}$. The ratio of $\tilde{p}_{t,\text{heuristic}}$ to $\tilde{p}_{t,\text{non-heuristic}}$ in the problem text is 1:8. In this experiment, we use a model that could produce reasonable output even when the context length increases. Specifically, we use only the GPT-3.5-turbo and GPT-4.
Figure 4: Number of cases where a particular distractor is selected (y-axis: $r$) in each reasoning step (x-axis: $d$). The heuristic to be included in the distructor is overlap.
Figure 5: Number of cases where a particular distractor is selected (y-axis: $r$) in each reasoning step (x-axis: $d$). The heuristic to be included in the distructor is position.
Figures 4 and 5 show how many times out of 100 times the model selected $\tilde{p}_{t,\text{heuristic}}$, $\tilde{p}_{t,\text{non-heuristic}}$, and $h^{*}_{t}$ at each step. Figures 4 and 5 show the experimental results for the cases where the heuristics included in the distructor are overlap and position, respectively. Figures 4 and 5 show that the number of cases in which the shortest path is selected increases as the goal is reached. The ratio of distructors that include heuristics to those that do not is 1:8, but the ratio of distructors that include heuristics when distructors are selected is higher than $1/(1+8)$. Therefore, this suggests that premises are being selected using heuristics. Except for the GPT-3.5 (overlap) results, the number of cases where a destructor is selected decreases as the goal is approached. These results are as with the results for 4.2. The results of GPT-3.5 (overlap) may indicate that the increase in the number of distructors has reduced the focus on distructors, including heuristics.
Answer the context question using the following example.
@@ Appendix H Effects of increasing the number of distructors
Context: Alice has 92 more bananas than Quentin’s mother. Victor has 10 fewer bananas than Walter. Xavier has 59 more bananas than Sybil. Yvonne has 79 more bananas than Sybil. Judy has 23 more bananas than Alice. Dave has 60 more bananas than Victor. Quentin has 35 fewer bananas than Quentin’s father. Heidi has 95 more bananas than Victor. Ursula doesn’t have 32 more bananas than Quentin’s father. Larry has 17 fewer bananas than Alice. Zoe has 58 fewer bananas than Yvonne. Ivan has 43 fewer bananas than Yvonne. Walter has 43 fewer bananas than Quentin’s mother. Nancy has 34 bananas. Grace has 41 more bananas than Xavier. Quentin’s mother has 55 fewer bananas than Nancy. Sybil has 3 fewer bananas than Nancy. Quentin’s father has 50 more bananas than Walter. Trent has 33 fewer bananas than Xavier.
@@
Nancy has 34 bananas, and Quentin’s mother has 55 fewer bananas than Nancy. So, Quentin’s mother has 34-55=-21 bananas.
Quentin’s mother has -21 bananas, and Walter has 43 fewer bananas than Quentin’s mother. So, Walter has -21-43=-64 bananas.
@@
Quentin’s father has -14 bananas, and Quentin has 35 fewer bananas than Quentin’s father. So, Quentin has -14-35=-49 bananas.
@@
Context: Yvonne’s father has 10 more apples than Yvonne’s son. Eve has 2 apples. Yvonne’s son has 3 more apples than Eve. Quentin has 3 more apples than Yvonne. Yvonne has 3 fewer apples than Yvonne’s father. Alice has 3 more apples than Grace. Trent has 34 more apples than Yvonne’s father. Ivan has 3 apples. Ursula has 3 more apples than Yvonne’s father. Grace has 3 apples. Xavier has 3 more apples than Ivan.
@@
Yvonne’s father has 15 apples, and Yvonne has 3 fewer apples than Yvonne’s father. So, Yvonne has 15-3=12 apples.
@@
Context: Kevin’s friend has 33 fewer grapes than Rob. Ivan has 43 more grapes than Victor. Victor has 33 fewer grapes than Kevin’s friend. Ursula has 75 fewer grapes than Zoe. Alice has 11 more grapes than Eve. Dave has 11 more grapes than Eve. Olivia has 29 more grapes than Kevin’s friend. Mallory has 97 more grapes than Olivia. Judy has 78 more grapes than Olivia. Rob has 55 grapes. Frank has 70 fewer grapes than Heidi. Eve has 84 fewer grapes than Kevin’s neighborhood. Xavier has 36 more grapes than Heidi. Kevin’s neighborhood has 55 fewer grapes than Kevin’s friend. Kevin has 43 fewer grapes than Kevin’s mother. Heidi has 61 fewer grapes than Kevin’s friend. Kevin’s mother has 88 more grapes than Kevin’s neighborhood. Kevin’s friend has 40 more grapes than Rob. Walter has 38 more grapes than Victor.
@@
Kevin’s friend has 95 grapes, and Kevin’s neighborhood has 55 fewer grapes than Kevin’s friend. So, Kevin’s neighborhood has 95-55=40 grapes.
@@
Kevin’s mother has 128 grapes, and Kevin has 43 fewer grapes than Kevin’s mother. So, Kevin has 128-43=85 grapes.
The final answer is 85.
@@
Answer the context question using the following example.
@@
Context: Nancy has 34 bananas. Mallory has 55 fewer bananas than Nancy. Walter has 43 fewer bananas than Mallory. Peggy has 50 more bananas than Walter. Quentin has 35 fewer bananas than Peggy. Alice has 92 more bananas than Mallory. Victor has 10 fewer bananas than Walter. Xavier has 59 more bananas than Sybil. Yvonne has 79 more bananas than Sybil. Judy has 23 more bananas than Alice. Dave has 60 more bananas than Victor. Heidi has 95 more bananas than Victor. Ursula doesn’t have 32 more bananas than Peggy. Larry has 17 fewer bananas than Alice. Zoe has 58 fewer bananas than Yvonne. Ivan has 43 fewer bananas than Yvonne. Grace has 41 more bananas than Xavier. Sybil has 3 fewer bananas than Nancy. Trent has 33 fewer bananas than Xavier.
@@
Nancy has 34 bananas, and Mallory has 55 fewer bananas than Nancy. So, Mallory has 34-55=-21 bananas.
Mallory has -21 bananas, and Walter has 43 fewer bananas than Mallory. So, Walter has -21-43=-64 bananas.
@@
Peggy has -14 bananas, and Quentin has 35 fewer bananas than Peggy. So, Quentin has -14-35=-49 bananas.
@@
Context: Eve has 2 apples. Yvonne’s son has 3 more apples than Eve. Zoe has 10 more apples than Yvonne’s son. Yvonne has 3 fewer apples than Zoe. Alice has 3 more apples than Grace. Quentin has 3 more apples than Yvonne. Trent has 34 more apples than Zoe. Ivan has 3 apples. Ursula has 3 more apples than Zoe. Grace has 3 apples. Xavier has 3 more apples than Ivan.
@@
Zoe has 15 apples, and Yvonne has 3 fewer apples than Zoe. So, Yvonne has 15-3=12 apples.
@@
Context: Rob has 55 grapes. Trent has 40 more grapes than Rob. Sybil has 55 fewer grapes than Trent. Zoe has 88 more grapes than Sybil. Kevin has 43 fewer grapes than Zoe. Kevin’s friend has 33 fewer grapes than Rob. Ivan has 43 more grapes than Victor. Victor has 33 fewer grapes than Kevin’s friend. Ursula has 75 fewer grapes than Zoe. Alice has 11 more grapes than Eve. Dave has 11 more grapes than Eve. Olivia has 29 more grapes than Kevin’s friend. Mallory has 97 more grapes than Olivia. Judy has 78 more grapes than Olivia. Frank has 70 fewer grapes than Heidi. Eve has 84 fewer grapes than Sybil. Xavier has 36 more grapes than Heidi. Heidi has 61 fewer grapes than Trent. Walter has 38 more grapes than Victor.
@@
Trent has 95 grapes, and Sybil has 55 fewer grapes than Trent. So, Sybil has 95-55=40 grapes.
@@
Zoe has 128 grapes, and Kevin has 43 fewer grapes than Zoe. So, Kevin has 128-43=85 grapes.
The final answer is 85.
@@
Answer the context question using the following example.
@@
Context: Alice has 92 more bananas than Mallory. Victor has 10 fewer bananas than Walter. Xavier has 59 more bananas than Sybil. Yvonne doesn’t have 79 more bananas than Sybil. Judy doesn’t have 23 more bananas than Alice. Dave has 60 more bananas than Victor. Quentin has 35 fewer bananas than Peggy. Heidi has 95 more bananas than Victor. Ursula doesn’t have 32 more bananas than Peggy. Larry doesn’t have 17 fewer bananas than Alice. Zoe has 58 fewer bananas than Yvonne. Ivan has 43 fewer bananas than Yvonne. Walter has 43 fewer bananas than Mallory. Nancy has 34 bananas. Grace doesn’t have 41 more bananas than Xavier. Mallory has 55 fewer bananas than Nancy. Sybil doesn’t have 3 fewer bananas than Nancy. Peggy has 50 more bananas than Walter. Trent doesn’t have 33 fewer bananas than Xavier.
@@
Nancy has 34 bananas, and Mallory has 55 fewer bananas than Nancy. So, Mallory has 34-55=-21 bananas.
Mallory has -21 bananas, and Walter has 43 fewer bananas than Mallory. So, Walter has -21-43=-64 bananas.
@@
Peggy has -14 bananas, and Quentin has 35 fewer bananas than Peggy. So, Quentin has -14-35=-49 bananas.
@@
Context: Zoe has 10 more apples than Yvonne’s son. Eve has 2 apples. Yvonne’s son has 3 more apples than Eve. Quentin has 3 more apples than Yvonne. Yvonne has 3 fewer apples than Zoe. Alice has 3 more apples than Grace. Trent has 34 more apples than Zoe. Ivan has 3 apples. Ursula has 3 more apples than Zoe. Grace has 3 apples. Xavier doesn’t have 3 more apples than Ivan.
@@
Zoe has 15 apples, and Yvonne has 3 fewer apples than Zoe. So, Yvonne has 15-3=12 apples.
@@
Context: Kevin’s friend has 33 fewer grapes than Rob. Ivan doesn’t have 43 more grapes than Victor. Victor doesn’t have 33 fewer grapes than Kevin’s friend. Ursula has 75 fewer grapes than Zoe. Alice has 11 more grapes than Eve. Dave has 11 more grapes than Eve. Olivia doesn’t have 29 more grapes than Kevin’s friend. Mallory has 97 more grapes than Olivia. Judy doesn’t have 78 more grapes than Olivia. Rob has 55 grapes. Frank has 70 fewer grapes than Heidi. Eve has 84 fewer grapes than Sybil. Xavier doesn’t have 36 more grapes than Heidi. Sybil has 55 fewer grapes than Trent. Kevin has 43 fewer grapes than Zoe. Heidi has 61 fewer grapes than Trent. Zoe has 88 more grapes than Sybil. Trent has 40 more grapes than Rob. Walter has 38 more grapes than Victor.
@@
Trent has 95 grapes, and Sybil has 55 fewer grapes than Trent. So, Sybil has 95-55=40 grapes.
@@
Zoe has 128 grapes, and Kevin has 43 fewer grapes than Zoe. So, Kevin has 128-43=85 grapes.
The final answer is 85.
@@
Appendix I Usage of Writing Assistance
We use publicly available writing assistance tools, including Grammarly, to refine the language for readability.